# Quantum Geometry Induced Kekulé Superconductivity in Haldane phases

### v1

@@ Abstract
Chiral Kekulé superconductivity.—The proposed superconducting state exhibits a spatial Kekulé pattern on the projected bipartite lattice, indicated in Fig. 4 c, and importantly, the size of the superconducting unit cell is tripled. The odd $J$-chiral Kekulé superconductor is a valley triplet with the spontaneously broken symmetry $U(1)\otimes O_{V}(3)\otimes\mathbb{Z}_{3}$, where the $U(1)$ is associated with the overall superconducting phase, $O_{V}(3)$ is the valley symmetry, and $\mathbb{Z}_{3}$ is related to the lattice translational symmetry. Alternatively, the even $J$-chiral Kekulé superconductor must be a valley singlet with the spontaneously broken symmetry $U(1)\otimes\mathbb{Z}_{3}$. Unlike the phase-modulated superconductors revealed in the quarter-metal regime [45, 42], the Kekulé superconductors in the half-metal regime are characterized by a spatial modulation of the amplitude of the superconducting order parameter (Fig. 4 c), which can be directly probed using scanning tunneling microscopy.
@@
Kekulé superconducting textures have previously been proposed for Dirac fermions in graphene, arising from next-nearest-neighbor pairing interactions [53, 63]. However, the mechanism we describe here is fundamentally different: it originates from topologically enforced geometric constraints of the Bloch wavefunctions, which enhances the pair susceptibility at large momenta ($\pm 2{\bf K}_{D}$). As such, this Kekulé superconducting order can emerge from a broad class of attractive interactions in the Haldane phase of C2DEGs, independent of the detailed microscopic pairing mechanism.
@@
Application to $J$-layer RG.—We now discuss the relevance of our findings to RG. In RG, trigonal warping of the bands, due to remote hopping processes, induces Lifshitz transitions [64], wherein the Fermi surface evolves from being circularly connected to an annular Fermi surface or a set of disconnected pockets. For tetralayer RG, this transition occurs in the half-metal regime at an electron density of $n_{e}\sim 0.4\times 10^{12}$ cm-2 near zero displacement field. Even above such a density, where the Fermi surface remains connected, it loses its azimuthal symmetry due to trigonal warping. In the presence of a trigonal warping energy scale $E_{tri}$, the weak-coupling enhancement of the intra-valley pair susceptibility is cut off by $E_{tri}$, and Eq. 3 retains the same functional form, but with $\log(E_{c}/T)\to\log(E_{c}/E_{tri})$. Consequently, a critical pairing strength $G_{+,-J}=-(1-x^{2})^{\alpha}/(N_{0}\log(E_{c}/E_{tri}))$ is required to stabilize Kekulé superconductivity. The general continuum Hamiltonian for a $J$-layer RG, including remote hopping, can be written as a sum of C2DEG Hamiltonians [65, 66], implying that the chiral index of the Kekulé superconductor can become density dependent. Nevertheless, because the Kekulé superconductivity results from the topologically enforced quantum geometry of the Bloch states, it should be robust against trigonal-warping effects so long as the Fermi surface remains connected.
@@
Discussions and Outlook.—In conventional metals, the emergence of PDWs typically requires strong correlation effects or fluctuations tied to lattice-scale physics, often leading to modulation periods incommensurate with the crystalline unit cell [69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86]. In contrast, we showed that in the Haldane phase of C2DEGs, the pair susceptibility is enhanced at finite momentum due to the topologically enforced quantum geometry of the Bloch wavefunctions. This enhancement occurs at high-symmetry points in the Brillouin zone, naturally favoring the formation of lattice-scale PDW states even in the weak-coupling regime. Unlike conventional scenarios that rely on strong-coupling mechanisms, competing orders, or fluctuation-driven instabilities, the quantum geometric mechanism proposed here offers a topologically protected mechanism to realize a lattice-scale PDW order in other chiral Chern bands of 2D crystals and layered materials.
@@ I Acknowledgements
Work by Y.B. was supported in part by the DOE EPSCoR program under the award DE-SC0022178, and a grant from NIST. F.Z. was supported by the National Science Foundation under grants DMR-2414726, DMR-1945351, and DMR-2324033 and by the Welch Foundation under grant AT-2264-20250403. Work by E.R. was funded by the US Department of Energy, Office of Basic Energy Sciences, via Award DE-SC0022245. Two (Y.B and E.R.) of us would like to recognize the Kavli Institute for Theoretical Physics (KITP) for hosting us as part of the moire24 program, which is supported by NSF PHY-2309135, where part of this research was performed.

### v2

@@ Abstract
Chiral Kekulé superconductivity.—The proposed superconducting state exhibits a spatial Kekulé pattern on the projected bipartite lattice, indicated in Fig. 4 d, and notably, the size of the superconducting unit cell is tripled. The odd $J$-chiral Kekulé superconductor is a valley triplet with the spontaneously broken symmetry $U(1)\otimes O_{V}(3)\otimes\mathbb{Z}_{3}$, where the $U(1)$ is associated with the overall superconducting phase, $O_{V}(3)$ is the valley symmetry, and $\mathbb{Z}_{3}$ is related to the lattice translational symmetry. Alternatively, the even $J$-chiral Kekulé superconductor must be a valley singlet with the spontaneously broken symmetry $U(1)\otimes\mathbb{Z}_{3}$. Unlike the phase-modulated superconductors revealed in the quarter-metal regime [45, 42], the Kekulé superconductors in the half-metal regime are characterized by a spatial modulation of the amplitude of the superconducting order parameter (Fig. 4 d), which can be directly probed using scanning tunneling microscopy.
@@
Kekulé superconducting textures have previously been proposed for Dirac fermions in graphene, arising from nearest-neighbor pairing interactions [54, 64]. However, the mechanism we describe here is fundamentally different: it originates from topologically enforced geometric constraints of the Bloch wavefunctions, which enhances the pair susceptibility at large momenta ($\pm 2{\bf K}_{D}$). As such, this Kekulé superconducting order can emerge from a broad class of attractive interactions in the Haldane phase of C2DEGs, independent of the detailed microscopic pairing mechanism.
@@
In RG, trigonal warping of the bands, due to remote hopping processes, induces Lifshitz transitions [65], wherein the Fermi surface evolves from being circularly connected to an annular Fermi surface or a set of disconnected pockets. For tetralayer RG, this transition occurs in the half-metal regime at an electron density of $n_{e}\sim 0.4\times 10^{12}$ cm-2 near zero displacement field. Even above such a density, where the Fermi surface remains connected, it loses its azimuthal symmetry due to trigonal warping. In the presence of a trigonal warping energy scale $E_{tri}$, the weak-coupling enhancement of the intra-valley pair susceptibility is cut off by $E_{tri}$, and Eq. 3 retains the same functional form, but with $\log(E_{c}/T)\to\log(E_{c}/E_{tri})$. Consequently, a critical pairing strength $G_{+,-J}=-(1-x^{2})^{\alpha}/(N_{0}\log(E_{c}/E_{tri}))$ is required to stabilize Kekulé superconductivity for even chirality. The general continuum Hamiltonian for a $J$-layer RG, including remote hopping, can be written as a sum of C2DEG Hamiltonians [66, 67], implying that the chiral index of the Kekulé superconductor can become density dependent. Nevertheless, because the Kekulé superconductivity results from the topologically enforced quantum geometry of the Bloch states, it should be robust against trigonal-warping effects so long as the Fermi surface remains connected.
@@
Discussions and Outlook.—In conventional metals, the emergence of PDWs typically requires strong correlation effects or fluctuations tied to lattice-scale physics, often leading to modulation periods incommensurate with the crystalline unit cell [70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87]. In contrast, we showed that in the Haldane phase of C2DEGs, the pair susceptibility is enhanced at finite momentum due to the topologically enforced quantum geometry of the Bloch wavefunctions. This enhancement occurs at high-symmetry points in the Brillouin zone, naturally favoring the formation of lattice-scale PDW states even in the weak-coupling regime. Unlike conventional scenarios that rely on strong-coupling mechanisms, competing orders, or fluctuation-driven instabilities, the quantum geometric mechanism proposed here provides a topologically protected mechanism to realize a lattice-scale PDW order in other chiral Chern bands of 2D crystals and layered materials.
@@ I Acknowledgements
Work by Y.B. was supported in part by the DOE EPSCoR program under the award DE-SC0022178, and a grant from NIST. F.Z. was supported by the National Science Foundation under grants DMR-2414726, DMR-1945351, and DMR-2324033 and by the Welch Foundation under grant AT-2264-20250403. Work by E.R. was funded by the US Department of Energy, Office of Basic Energy Sciences, via Award DE-SC0022245. Two of us (Y.B. and E.R.) would like to recognize the Kavli Institute for Theoretical Physics (KITP) for hosting us as part of the moire24 program, which is supported by NSF PHY-2309135, where part of this research was performed.